Find the equation ,in the form y = ax^2 + bx + c,of the parabola which passes through the points(-3,0) and (1,-16).
Found 2 solutions by Alomelo, Alan3354:
Answer by Alomelo(2)

For (-3,0)
we write 0=(-3)^2+b(-3)+C, So, 3b-c=9........(i)
and for(1,-16), So b+c=-17......................(ii)

So (i)+(ii) we get b=-2
So put this into equation(ii) we get c=-15
so ultimetly the equation is: y=x^2-2x-15.

Answer by Alan3354(69443)

Find the equation ,in the form y = ax^2 + bx + c,of the parabola which passes through the points(-3,0) and (1,-16).
--------------
2 points is not sufficient to define a parabola, 3 are needed.
An infinite number of parabolas can be found that pass thru 2 points.